50 recent changes in AdaptiveOptics Web retrieved at 02:23 (GMT)

20220514-22_LBTI_Night_TN
Compare Flux SX DX Selected tracking numbers taken nearly simultaneously DX TrackNo DXMag DXRec DXbin DX#mod DXfreq DXemGain DXgain DXmod DXnp...
20220524_SX_Day
Details. Power on. * OK camera lens * not live. * Power cycle bcu2k restore setup ok * frames:20220524_155800 * one quadrant more illuminated. *...
Logs
AO Daily Logs May 2022. 20220524_SX_Day . Close loop check. Pupil wobble 20220514 22_LBTI_Night_TN . Tracking Number Tables for LBTI May run 20220516_SX_Da...
AOSWTest-20220425-idlsystem
2022 04 25 verify AbstractSystem::_cmdResult thread safety bug SW: BMAO: AV, JCG Software testing queue Servers and software versions used: sxadsec: stablelbti ...
20220516_SX_Day
Details. Power on OK Ocam OK. test preset. Mon May 16 18:00:05.751 2022 aos.PresetAO{FLAO}.started PresetAO {FLAO} 2236 (aomode:ACE AO,focst:bentGregorianFront...
20220515_DX_Day
20220515_DX_Day. DX ASM 20:23 (UT) testing magnets 20:31 (UT) nominal flat applied with 554 modes limiting actuator 440 (in the first ring from the center), but ...
20220514_SX_Night
1:10 Power up SX LUCI SOUL WFS 7:51 Adsec Failure at end of LBTI operations MD TIMEOUT CA * MirrorCtrl _106542 ERR 773 2022 05 14 07:51:15.223768 BCU...
AOSoftwareTests
AO Software Test Plans Upcoming/drafts * adsec resumeao gain ramp: test gain ramp failures Past * 2022 05 10 adsec switch bcu communication proble...
AOSWTest-20220510-setz0
2022 05 10 switch bcu communication problem: test changing to which banks set_z0.pro writes (re do with FAKESLOPE enabled and fixed IDL procedures) * Softwar...
AOSWTest-20220507-08-lbti-luci-checkouts
2022 05 07/08 (adsec soul sx wfs lbti sx wfs) release checkout: upgrade on LBTI and SOUL LUCI DEC nights * Software: Brandon: I'll have notification volu...
20220513_DX_Night
20220513_DX_Night. LBTI AO, Day 1 07:00 (UT) change configuration * cp bad_actuators.list.mmc bad_actuators.list * run idl in the directory $ADOPT_ROOT/conf...
20220513_SX_Night
LBTB 02:25 * OSCO First. * housekeeper.L INF 729146 2022 05 13 02:58:43.329267 SwitchSafeSkipFrameCnt 6444 housekeeper.L INF 729154 2022 05...
20220512_SX_Night
LBTB 02:41 * Starting with ESM. * Bin2 400 Hz 600 EM gain * mag 14.2 * g_tt_3.54_ho1_9.14_ho2_0.08.fits 03:18 * New target. * Looking for one that...
20220511_DX_Day
20220511_DX_Day. LBTI AO, Day 5 16:00 UT Changing ASM configuration to daytime 16:05 UT set flat 565 (force 0.04) RB temperature 5.8 C 20:32 UT Changing ASM conf...
20220511_DX_Day_2
20220510_DX_Day_2 DX S LUCI Set_Z0 test, Day 2 9:45 Starting * w_stop OK * w_check Ok * w_start Housekeeper did not start * w_stop OK * w_check OK ...
20220511_SX_Night
LBTB 02:07 * Opening, ready for OSCO. * Stable WITH NCPA. Every once in a while, we see a dark bit at the 10:00 position in the pupils. * Looks like we mig...
20220510_DX_Day_2
20220510_DX_Day_2. DX S LUCI Set_Z0 test, Day 1 17:30 Close loop with AOS Command GUI * In Center Pupils there are modes above TT that are corrected * LN...
20220510_DX_Day
20220510_DX_Day. LBTI AO, Day 4 15:04 UT Changing ASM configuration to daytime 15:07 UT Fastlink test failed on first program loading attempt, successful the seco...
20220509_SX_Night
LBTB Waiting for sunset. Using the new AO supervisor. TOO WINDY! 03:42 * The wind is calming down. * Starting with ESM. * Mag 12.4 * Optimize gain in ...
20220508_SX_Night
Details. AOeng@soul sxwfs source $ use_soulUsage: use_soul version Currently running: stable (soul)Available versions in /home/AOeng/soul: latest stable stable....
20220506_SX_Day
Details. 12:29 Power on * ok. Ocam * Triangle feature * power cycle bcu2k set ocam parameters(1x1 cropped). * ok * tn:_20220506_193234 12:49 powe ...
20220502_SX_Day
Details. Power on 21:29:49.833828 MAIN Cannot turn on PROG. Reason: Timeout waiting for powerboard.L.PROG_BCU_2K.CUR TIMEOUT_ERROR GENERIC_SYS_ERRORWfsArbi...
20220505_DX_Day
20220505_DX_Day. LBTI AO, Day 3 15:13 UT current chamber temperature 4.3C 15:53: UT current chamber temperature 4.8C, waiting for confirmation from monutain 15:59...
20220503_DX_Day
20220503_DX_Day. LBTI AO, Day 1 17:00 (UT) /local/aomeas/adsec_calib/flat/20220422_191142/flattening.sav has a small bump at 565 (force 0.12) * 17:00 (UT) R...
20220504_DX_Day
20220504_DX_Day. LBTI AO, Day 2 15:30 (UT) /local/aomeas/adsec_calib/flat/20220422_191142/flattening.sav has a small bump at 565 (force 0.26) 15:44 (UT) bumpin...
20220504_SX_Day
Summary. Details. Power up. ocam not live. pupil rotator find homes at 180 restart binning from ocam hw gui No success Power cycle bcu2k configure bin OK. Save ...
Procedures
Remove Bad or jump actuator. Remove Bad or jump actuator old notes . Clear daytime activity. Flash BCU47 firmware. Deployment Soul version environments. Cal...
20220503_SX_Day
Details. power on the system has the new power on file loaded. set flat. Adsec IDL controllwer down. start the process. adsc_stop/start. flat#635 test preset. OK ...
2022-03-pause-resume-loop-data
YYYY MM DD test gain ramp failure Software: BMAO: Software versions used: stable, test.YYYYMMDD.reorder gain rampSoftware version after test: stable Test 1: Ga...
20220501_DX_Day
20220501_DX_Day DX AdSec Calculate Influence Functions, Day 12 08:00 Wrote IDL routine to check if shape removal was included in the 4D processing of the phase ...
20220502_DX_Day
20220502_DX_Day. DX AdSec Calculate Influence Functions, Day 13 09:30 Perform IF analysis using routines from 2019 * copy current flao_environment.sh to ~/doug...
20220430_DX_Day
20220430_DX_Day DX AdSec Calculate Influence Functions, Day 11 Processed Set_4 from 20220425 where shapes (TT, Focus, Coma) were NOT removed in the 4D processing ...
20220425_DX_Day
20220425_DX_Day DX AdSec Influence Function Data Collection with New 4D, Day 9 3:30 start. Amali and SX LBTI in parallel 03:40 Processing of yesterday's IF data...
AOSWTest-20220425-uninterruptible
2022 04 25 make AO arbitrator commands, including OffsetXY, "uninterruptible" Software: BMAO: AV, JCG Software testing queue Servers and software versions used:...
AO_tests
Monitoring. * Pupil Wobble. * NCPA Tests. * BIN3 optimize Gopt. : To be scheduled * Misalignment and resume loop. * Possible reacquire a n...
20220427_DX_Day
20220427_DX_Day DX AdSec Calculate Influence Functions, Day 10 Set 1 from 20220424 11:50 Commands (following 20190908_DX_Day TWiki) * Tracking numbers IDL TN...
20220427_DX_Day_LBTI_Bad_Frames
DX LBTI Bad Frames Test 1 Pixels are all zero 2 Slopes are all zero 4 Some pixels are NaN 8 Some slopes are NaN 16 The RMS of the pixels is higher tha...
AdSec_Ele_Change
The AdSec software has been recently modified to allow changing the elevation limit for the AdSec via command line, without the need for a restart of the AdSec so...
2022-04-27Change_elevation_limitScriptUpgrade
(adsec) change_elevation_limit script upgrade test Software: XZAO: XZ, GB Problem DX ASM had a failure on April 26 08:55 UT. The cause is that the changing of el...
20220424_DX_Day
20220424_DX_Day DX AdSec IF data collection, Day 8 11:30 Arrive at LBTO Restart DX AdSec in MMC configuration 08:14 change FF matrix * cd /home/aoeng/soul/mi...
20220426_SX_Night
DEC/EDC/EIEIO 08:18 * Handover is at 0900 Zulu. * Seeing has turned to crap. * It was great a couple of hours ago. :( 09:19 * OSCO and the seeing has i...
20220424_SX_Night
Details. Power on 5:30PM * ocam Ok * osco. * sky hazy, most probably due the fires in Flagstaff. 7:04 PM collimation * el 50 7:27 OSCO * luci mos e...
AOSWTest-20220419-agwpower
2022 04 19 bcu2k setup timeouts: test agw_power fix that waits for RelayCtrl to be in OPERATING state before issuing power on commands * Time: daytime * S...
AOSWTest-upcoming-bcu2ktimeouts
2022 04 19 bcu2k setup timeouts: test agw_power fix that waits for RelayCtrl to be in OPERATING state before issuing power on commands * Time: daytime * S...
AOSWTest-20220425-copleyctrl
2022 04 25 bayside out of range movements: test that out of range bayside movements report errors immediately instead of waiting to time out * Software: BM ...
AOSWTest-20220425-optloopdata
2022 04 25 optical loop diagnostic data during pause, uninterruptible offsets, bayside out of range movement timeouts, verify idlsystem thread safety bug Softwa...
AOSWTest-20220425
2022 04 25 optical loop diagnostic data during pause, uninterruptible offsets, bayside out of range movement timeouts, verify idlsystem thread safety bug Softwa...
2022-04-swap-pause-order
upcoming verify AbstractSystem::_cmdResult thread safety bug SW: BMAO: Software testing queue Servers and software versions used: sxadsec: stable (no change)soul...
20220423_SX_Night
Details. Lost internet connection to the mountain. possible no AO. Dougn call JUan CArlos, internet is back to the mountain. Prep for AO night. 7:28PM power up. o...
20220422_DX_Day
20220422_DX_Day DX AdSec Flattening with New 4D, Day 7 Started at 6:45. RR on early 08:15 4D, DX AdSec aligned but pupil shifted 09:46 Pupil aligned 09:49 fsm_l...
Number of topics: 50
Page 1 of 26 Next >

See also: rss-small RSS feed, recent changes with 50, 100, 200, 500, 1000 topics, all changes
Topic revision: r1 - 26 Feb 2018, ProjectContributor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback